RC meeting and Run Plan, Thursday November 24th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information


RC Updates


  • For several hours, Bill under Dave Hamlette supervision, cleaned most of the target enclosure, but more work is still needed. After that, Bill and Arun replaced the damaged foils with new foils.
  • Jiwan with Simona help (Bill was sent home after a long day) tried to set up the beam on target. A discrepancy between what was assumed the carbon hole position in place, with what was measured with the daq, discouraged the use of the reference cell. The discussion was postponed to the next day morning (today).
  • With the carbon foils restored, and the beam restored into the hall (besides its use to the reference cell), Holly and Xinzhan proceed with the BB Gem luminosity scan up to 45uA.
  • Early this morning, Bill cross-checked the position of the carbon hole wrt the laser and found that the position is the same as before the target incident --> https://logbooks.jlab.org/entry/4090304
  • Since there is not issues with the position of the carbon hole, thus the position of the reference cell wrt the carbon hole, Bill. Jiwan and Mark started setting up the beam (position and width) and at the moment of this summary, they are ready to test to the empty cell and continue with the run plan discussed yesterday.
  • The two hodoscope channels were identifued by David Hamilton via Slack upon request from Maria:

30R has had an intermittent fault since GMn. We suspect poor optical contact due to a mechanical issue. Same problem for 2R, which appeared a few weeks back. 84R dropped when I was on shift on Friday. That channel is over-current so it’s a pmt/base issue. The top 10 bars are out of acceptance, so we are not too concerned. I suspect it will be after GEn before we get a good look at these issues.

Short term run plan

  • Integration of SBS GEMS to Main DAQ (who will be responsible and when?)
  • Take data on H2 reference cell
  • Take run with lower BBCAL threshold (~1 GeV) to accept pions
    • DAC must be changed from 366 to 265mV
    • According to Bogdan, start with H2, check rates and then change to N2.
    • As far I understood, Bogdan suggested to go first single arm and then coincidence. But not sure coincidence woth H2 included or just N2.
    • without more input, I would say one run per set-up, maybe 2h run? 2M events? Depending of the data rate.
  • When do we want the pion rate test? Today as soon we can or tomorrow Friday with more availability of the PIs?
  • Sieve slit installation on Saturday and then optics.
  • Prepare to install another cell (target group)


  • Shift schedule unfilled for next week
  • Dump right Ion chamber cabling needs to be changed (Jerry Kowal is aware of it)

Hall Access

  • No access programmed.

Summary of GEn data taking

As soon I get where to get the info, I will update this part.